OpenStudy (anonymous):

Find the surface area of a rectangular prism that has the following dimensions. length: 2 ft width: 14 in. height: 11 in.

jimthompson5910 (jim_thompson5910):

SA = 2*(LW + LH + WH) SA = 2*(24*14 + 24*11 + 14*11) SA = 2*(336 + 264 + 154) SA = 2*(754) SA = 1508 So the surface area of the rectangular prism is 1508 square inches.

jimthompson5910 (jim_thompson5910):

Keep in mind that you have to convert everything to the same unit of measure. So that's why I converted the length of 2 ft to 2*12 = 24 inches.